# Do Whitepaper Claims Predict Market Behavior? Evidence from Cryptocurrency Factor Analysis

### v1

@@ 6.3 DeFi Divergence
Application-layer tokens (ENS, UNI, AAVE) show the largest narrative-market gaps (0.014-0.014 to 0.035-0.035). ENS’s extreme divergence (0.035-0.035) reflects the disconnect between naming service utility and market dynamics. This pattern suggests a structural distinction: infrastructure tokens (NEAR, ATOM) may be priced on technical specifications—sharding, throughput, interoperability—the very dimensions whitepapers emphasize, while application tokens (UNI, ENS, AAVE) may be priced on usage metrics (TVL, fee revenue, governance participation) that whitepapers cannot anticipate. We note that the magnitude of these entity-level impacts (0.007–0.035) remains modest relative to overall alignment levels, so this interpretation should be treated as exploratory rather than definitive.
@@ 6.4 Feature Importance Patterns
Scalability and Governance claims contribute most to alignment (+0.014+0.014 and +0.010+0.010 respectively), while Security claims show the largest negative impact (0.022-0.022). This counterintuitive finding admits multiple interpretations: security claims may be ubiquitous across projects (most whitepapers emphasize cryptographic security) providing little discriminative signal, actual security-related market events (hacks, exploits) may create idiosyncratic volatility orthogonal to stated claims, or the BART-MNLI classifier may conflate security-specific terminology with general technical descriptions. Projects emphasizing scalability and governance differentiation show better narrative-market correspondence, perhaps because these claims map more directly to observable protocol characteristics (throughput metrics, voting mechanisms) that markets can assess.
@@ 6.5 Theoretical Implications
Narrative Dissociation Hypothesis. The weak alignment we document is consistent with what we term “narrative dissociation”—an observed weak correspondence between stated project intentions and realized market behavior. We emphasize that our limited sample size ($n=23$) provides insufficient power to definitively distinguish weak alignment from no alignment; we can confidently reject strong alignment ($\phi\geq 0.70$), but this framing represents a working hypothesis rather than a demonstrated finding. If genuine, narrative dissociation would contrast with efficient market theory, which predicts that informative narratives are rapidly incorporated into prices. The evolving dependency structures in cryptocurrency markets (Briola and Aste, 2022) and the documented role of social media in price dynamics (Burnie et al., 2020) suggest narrative-market relationships may be more complex than our static alignment tests capture. Our findings suggest that either narratives contain little price-relevant information, or markets systematically ignore such information—though we cannot adjudicate between these explanations with present data.
@@ 6.7 Limitations
  • Small sample size ($n=23$) is a critical constraint. With only 23 common entities, statistical power is limited to detecting strong alignment ($\phi\geq 0.70$); we cannot reliably distinguish weak alignment from no alignment. While our expanded corpus (24 whitepapers, 23 common entities) substantially improves over prior work, expanding to 50+ projects would enable both adequate power for moderate effects and subsample analysis by sector.
@@
  • Zero-shot classifiers trained on general-domain NLI corpora exhibit domain shift when applied to specialized cryptocurrency discourse (Gururangan et al., 2020). While BART-MNLI captures broad semantic categories, crypto-specific terminology (“sharding,” “AMM,” “tokenomics”) may not receive accurate treatment. Beyond acknowledging this limitation, we interpret it as a substantive methodological finding: the “Semantic Gap” between general-purpose NLP and crypto-native discourse represents a measurement challenge that future researchers must address. Our security ablation result—where security claims hurt alignment (0.022-0.022) despite security being technically salient—provides empirical evidence of this gap, likely reflecting BART-MNLI conflating cryptographic security (“zero-knowledge proofs,” “trusted execution environments”) with general safety concepts. This finding carries practical implications: off-the-shelf LLMs should not be deployed for cryptocurrency auditing or regulatory classification without domain adaptation via continued pretraining on cryptocurrency corpora.
@@ 7 Conclusions
We investigated whether cryptocurrency whitepaper claims predict market behavior using a novel pipeline combining NLP claims extraction, tensor decomposition, and Procrustes alignment. Our analysis yields nuanced results: while narrative-market alignment remains weak ($\phi<0.35$), the statistics–factors comparison achieves statistical significance ($p<0.001$), suggesting market metrics systematically relate to tensor-derived factors even as narrative claims remain decoupled.
@@ 7.1 Summary of Findings
Our investigation across 24 whitepapers and 23 common entities produced the following key findings:
@@
    Weak Narrative Alignment. Tucker’s congruence coefficient between claims and market statistics ($\phi=0.341$, $p=0.332$), and claims and tensor factors ($\phi=0.077$, $p=0.747$), both fall well below the 0.65 threshold for moderate similarity and fail to achieve statistical significance.
@@
    Significant Statistics–Factors Link. The statistics–factors comparison ($\phi=0.197$, $p<0.001$) achieves statistical significance despite modest magnitude, indicating market summary statistics systematically relate to latent factor structure.
@@
    Infrastructure Token Alignment. NEAR, MKR, and ATOM show modest positive alignment contributions (+0.007 to +0.009), while major assets (BTC, ETH) and DeFi tokens (ENS, UNI, AAVE) hurt alignment.
@@
    Bitcoin Robustness. Excluding Bitcoin ($n=22$) yields $\phi=0.081$ for claims–factors, confirming the null result is not driven by Bitcoin’s dominant market position.
@@
  6. 6.
    Institutional Corpus. Using a February 2025 institutional corpus (official documentation, foundation updates, governance posts, project blogs; 12 assets with complete market data), alignment remains weak ($\phi=0.203\pm 0.013$), indicating that broader narratives beyond whitepapers do not materially improve alignment.
@@ Appendix E Per-Dimension Alignment Values
Table 10: Per-Dimension Congruence Coefficients
Dim Claims–Stats Claims–Factors Stats–Factors
1 0.342 0.309 0.690
2 0.397 0.310 0.693
3 0.345 0.000 0.000
4 0.454 0.000 0.000
5 0.529 0.000 0.000
6 0.302 0.000 0.000
7 0.360 0.000 0.000
8 0.285 0.000 0.000
9 0.318 0.000 0.000
10 0.000 0.000 0.000
Mean (all) 0.333 0.062 0.138
Mean (non-zero) 0.370 0.310 0.691
Notably, the statistics–factors comparison shows moderate alignment ($\phi\approx 0.69$) within the two non-padded dimensions, suggesting genuine correspondence between market statistics and tensor factors that is obscured by the mean across all dimensions. Claims-based comparisons remain limited even within non-padded dimensions ($\phi\approx 0.31$ for claims–factors).

### v2

@@ 6.3 DeFi Divergence
DeFi infrastructure tokens (RPL, HBAR, AAVE, SUSHI) show the largest negative contributions (0.011-0.011 to 0.018-0.018). RPL’s negative impact (0.018-0.018) reflects the disconnect between staking infrastructure claims and market dynamics. This pattern suggests a structural distinction: specialized tokens with clear market niches (XMR for privacy, CRV/YFI for yield) may be priced on the specific functionality their whitepapers describe, while broad DeFi infrastructure tokens may be priced on network effects and liquidity metrics that whitepapers cannot anticipate. We note that the magnitude of these entity-level impacts ($\pm$0.020) remains modest relative to overall alignment levels, so this interpretation should be treated as exploratory rather than definitive.
@@ 6.4 Feature Importance Patterns
Monetary claims (medium_of_exchange) contribute most to alignment (+0.021+0.021), followed by interoperability (+0.017+0.017) and privacy (+0.017+0.017). This pattern suggests markets reward projects with clear transactional and cross-chain value propositions. Data storage shows the only negative impact (0.001-0.001), though minimal. Projects emphasizing core monetary and infrastructure functionality exhibit better narrative-market correspondence than those with diffuse technical claims.
@@ 6.5 Theoretical Implications
Narrative Dissociation Hypothesis. The weak alignment we document is consistent with what we term “narrative dissociation”—an observed weak correspondence between stated project intentions and realized market behavior. We emphasize that our limited sample size ($n=37$) provides insufficient power to definitively distinguish weak alignment from no alignment; we can confidently reject strong alignment ($\phi\geq 0.70$), but this framing represents a working hypothesis rather than a demonstrated finding. If genuine, narrative dissociation would contrast with efficient market theory, which predicts that informative narratives are rapidly incorporated into prices. The evolving dependency structures in cryptocurrency markets (Briola and Aste, 2022) and the documented role of social media in price dynamics (Burnie et al., 2020) suggest narrative-market relationships may be more complex than our static alignment tests capture. Our findings suggest that either narratives contain little price-relevant information, or markets systematically ignore such information—though we cannot adjudicate between these explanations with present data.
@@ 6.7 Limitations
  • Small sample size ($n=37$) is a critical constraint. With only 37 common entities, statistical power is limited to detecting strong alignment ($\phi\geq 0.70$); we cannot reliably distinguish weak alignment from no alignment. While our expanded corpus (38 whitepapers, 37 common entities) substantially improves over prior work, expanding to 50+ projects would enable both adequate power for moderate effects and subsample analysis by sector.
@@
  • Zero-shot classifiers trained on general-domain NLI corpora exhibit domain shift when applied to specialized cryptocurrency discourse (Gururangan et al., 2020). While BART-MNLI captures broad semantic categories, crypto-specific terminology (“sharding,” “AMM,” “tokenomics”) may not receive accurate treatment. Beyond acknowledging this limitation, we interpret it as a substantive methodological finding: the “Semantic Gap” between general-purpose NLP and crypto-native discourse represents a measurement challenge that future researchers must address. This finding carries practical implications: off-the-shelf LLMs should not be deployed for cryptocurrency auditing or regulatory classification without domain adaptation via continued pretraining on cryptocurrency corpora.
@@ 7 Conclusions
We investigated whether cryptocurrency whitepaper claims predict market behavior using a novel pipeline combining NLP claims extraction, tensor decomposition, and Procrustes alignment. Our analysis yields nuanced results: while narrative-market alignment remains weak ($\phi<0.25$), the statistics–factors comparison achieves statistical significance ($p<0.001$), suggesting market metrics systematically relate to tensor-derived factors even as narrative claims remain decoupled.
@@ 7.1 Summary of Findings
Our investigation across 38 whitepapers and 37 common entities produced the following key findings:
@@
    Weak Narrative Alignment. Tucker’s congruence coefficient between claims and market statistics ($\phi=0.246$), and claims and tensor factors ($\phi=0.058$), both fall well below the 0.65 threshold for moderate similarity.
@@
    Significant Statistics–Factors Link. The statistics–factors comparison ($\phi=0.174$, $p<0.001$) achieves statistical significance despite weak magnitude, indicating market summary statistics systematically relate to latent factor structure.
@@
    Specialized Token Alignment. XMR, CRV, YFI, and SOL show positive alignment contributions (+0.010 to +0.020), while DeFi infrastructure tokens (SUSHI, AAVE, HBAR, RPL) hurt alignment.
@@
    Temporal Dynamics. Alignment shows moderate variation across six temporal windows ($\phi=0.162\pm 0.026$), ranging from $\phi=0.138$ to $\phi=0.200$, reflecting market regime changes.
@@
    Decomposition Robustness. Both CP and Tucker decomposition yield similar results ( 92.5% explained variance, comparable $\phi\approx 0.06$), ruling out method-specific artifacts.
  6. 6.
@@ Appendix E Per-Dimension Alignment Values
Table 14: Per-Dimension Alignment Coefficients (Claims–Statistics)
Dim Category $\phi$ Interpretation
1 store_of_value 0.241 Weak
2 medium_of_exchange 0.370 Weak
3 smart_contracts 0.181 Weak
4 defi 0.400 Weak
5 governance 0.577 Weak
6 scalability 0.445 Weak
7 privacy 0.242 Weak
8–10 (padding) 0.000 N/A
Mean 0.246 Weak
All dimensions show weak alignment ($\phi<0.65$), with governance showing the highest individual dimension alignment ($\phi=0.577$). Even the best-aligned claim category fails to reach moderate similarity, confirming the weak narrative-market correspondence is not driven by averaging across poorly-aligned categories.
@@ Appendix F Methodological Extensions for Future Work
Appendix G Per-Category Method Agreement
Figure 9 visualizes pairwise method correlations for each semantic category. While most categories exhibit positive inter-method agreement (r=0.4r=0.40.80.8), smart_contracts shows anomalous negative correlation between BART-NLI and embeddings (r=0.17r=-0.17), suggesting this category’s linguistic markers are interpreted differently across model architectures. Categories with clearer linguistic anchors (medium_of_exchange, DeFi, governance) show strongest convergence.
Figure 9: Per-category Pearson correlations between three classification methods: BART-NLI vs Embedding, BART-NLI vs LLM, and Embedding vs LLM. Most categories show moderate positive agreement (green), but smart_contracts exhibits negative BART–Embedding correlation (red), indicating model-specific interpretation of this technically ambiguous category.